# Multiple Imputation for Nonresponse in Complex Surveys Using Design Weights and Auxiliary Margins

### v1

@@ 4 Simulation Studies
To demonstrate the methodology, we mimic the simulation study used by Yang and Reiter (2025). They found that, by leveraging auxiliary margins, multiple imputation via the hybrid missingness MD-AM model can offer more accurate inferences than multiple imputation that does not utilize $\mathcal{A}$. However, they presume design weights are unused or unavailable for unit nonrespondents, whereas we use them in the imputation procedure.
We generate a population $\mathcal{P}$ comprising $N=3,373,378$ units. This is the number of individuals in the 2022 American Community Survey data used by Yang and Reiter (2025). We make one of the variables in this file a size variable $Z=(z_{1},\dots,z_{N})$; this is observed for all $N$ individuals. To generate a $\mathcal{S}$, we sample each record from $\mathcal{P}$ independently with $\pi_{i}=1/10z_{i}$, where $i=1,\dots,N$. This Poisson sampling design results in approximately $n=6000$ units in any $\mathcal{S}$, although the precise number varies. Let $W=10Z$ be the survey weights.
Following Yang and Reiter (2025), for each unit $i=1,\dots,N$, we generate four binary variables $(x_{i1},x_{i2},x_{i3},x_{i4})$ and two continuous variables $(x_{i5},x_{i6})$. We make $X_{1}$ depend on $W$ to ensure associations between the weights and the survey variables. The data generation model is displayed in (4.1)–(4.5).
@@
We independently sample 500 datasets using Poisson sampling. For each sample $\mathcal{S}$, we generate the unit nonrespondents by sampling Bernoulli random draws of $U_{i}$, where $i\in\mathcal{S}$, from the distribution of $U|X_{1},X_{2}$ implied by (4.1)–(4.3). We make $\boldsymbol{x}_{i}$ completely missing for any unit in $\mathcal{S}$ where the drawn $U_{i}=1$. For item nonresponse, we use the ICIN mechanisms defined in (4.6) for $X_{j}$ where $j\in\{2,3,4,6\}$, and we let $X_{1}$ and $X_{5}$ be fully observed.
@@
We presume that $T_{1}=\sum_{i=1}^{N}x_{i1}$ and $T_{2}=\sum_{i=1}^{N}x_{i2}$ are known and use them as $\mathcal{A}$. We use a bespoke implementation of the “mice” package in R (Van Buuren, 2018) to generate $L=10$ completed data sets for item nonresponse. As values for $V_{1}$ and $V_{2}$, we use MICE to make one completed dataset for all of $\mathcal{S}$ and compute the expressions for the usual design-based estimators of the variances of $\hat{T}_{1}$ and $\hat{T}_{2}$ under Poisson sampling. We create donor pools for the hot deck per Section 3.3 by matching on $(X_{1},X_{2})$.
After generating $L=10$ completed datasets, we use the inferential methods in Rubin (1987) for the estimands in Yang and Reiter (2025), including marginal totals for $(X_{1},\dots,X_{6})$ and several probabilities involving multiple categorical variables. In each completed dataset, we compute Horvitz and Thompson (1952) point and variance estimators for a Poisson sampling design based on the design weights for all sampled records. The results will be forthcoming in a revised version of this manuscript.

### v3

@@ 4 Simulation Studies
To demonstrate the methodology, we mimic the simulation study in Yang and Reiter (2025). They found that, by leveraging auxiliary margins, multiple imputation via the hybrid missingness MD-AM model can offer more accurate inferences than multiple imputation that does not utilize $\mathcal{A}$. However, they presume design weights are unused or unavailable for unit nonrespondents, whereas we use them per Section 3.
We generate $\mathcal{P}$ comprising $N=3,373,378$ units. This is the number of individuals in the 2022 American Community Survey data file used by Yang and Reiter (2025). We make one of the variables in this file a size variable $Z=(z_{1},\dots,z_{N})$; this is observed for all $N$ individuals. To generate a $\mathcal{S}$, we sample each record from $\mathcal{P}$ independently with $\pi_{i}=1/10z_{i}$, where $i=1,\dots,N$. This Poisson sampling design results in approximately $n=6000$ units in any $\mathcal{S}$. Let $W=10Z$ be the survey weights. For each unit $i=1,\dots,N$, we generate four binary variables $(x_{i1},x_{i2},x_{i3},x_{i4})$ and two continuous variables $(x_{i5},x_{i6})$. The data generation model is displayed in (4.1)–(4.5).
@@
We independently sample 500 datasets. For each sample $\mathcal{S}$, we generate the unit nonrespondents by sampling Bernoulli random draws of $U_{i}$, where $i\in\mathcal{S}$, from the distribution of $U|X_{1},X_{2}$ implied by (4.1)–(4.3). We make $\boldsymbol{x}_{i}$ completely missing for any unit in $\mathcal{S}$ where the drawn $U_{i}=1$. For item nonresponse, we use the ICIN mechanisms defined in (4.6) for $X_{j}$ where $j\in\{2,3,4,6\}$, and we let $X_{1}$ and $X_{5}$ be fully observed.
@@
We let $\mathcal{A}$ include $T_{1}=\sum_{i=1}^{N}x_{i1}$ and $T_{2}=\sum_{i=1}^{N}x_{i2}$, which we use for multiple imputation via the MD-AM models of Section 3. For Section 3.1, we generate $L=10$ completed datasets for item nonresponse using a bespoke implementation of the “mice” package in R. For Section 3.2, we consider two sets of working probabilities for $X_{1}$. The first allows $p_{i1c}$ to vary across units: each $p_{i1c}$ equals the unit’s predicted probability from the logistic regression of $X_{1}$ on $W$ estimated using the units with $U_{i}=0$. The second fixes $p_{i1c}=p_{1c}$ across units and uses (3.2). For both we use (3.9) and (3.10) to impute $X_{2}$. For the variances $V_{1}$ and $V_{2}$ used to sample $\hat{T}_{1}\sim N(T_{1},V_{1})$ and $\hat{T}_{2}\sim N(T_{2},V_{2})$, we use “mice” to make one completed dataset for all of $\mathcal{S}$ and compute the expressions for the unbiased variance estimators for $\hat{T}_{1}$ and $\hat{T}_{2}$ under Poisson sampling. For Section 3.3, we create donor pools for the hot deck by matching on $(X_{1},X_{2})$.
After generating $L=10$ completed datasets, we use the inferential methods in Rubin (1987) for the estimands in Yang and Reiter (2025), including marginal totals for $(X_{1},\dots,X_{6})$ and several probabilities involving multiple categorical variables. In each completed dataset, we compute Horvitz and Thompson (1952) point and variance estimators for a Poisson sampling design based on the design weights for all sampled records.
Figure XX summarizes the results for $\theta_{1}=-2$; results for $\theta_{1}=-0.5$ are available at URL HERE. Both MD-AM models produce accurate point estimates, mitigating the substantial bias from the nonignorable nonresponse. The multiple imputation confidence intervals have coverage rates exceeding 95% due to over-estimation of the variances. As discussed by Yang and Reiter (2025), this is because the combining rules of Rubin (1987) do not account for using known totals in imputation. Overall, the simulation results suggest that the MD-AM models from Section 3 offer principled ways to incorporate known marginal totals in multiple imputation when using the design weights for all sampled units.